OpenStudy (anonymous):

What is (x^2-3x+2) + (5x^2-3x-8)? A. -4x^2 + 10 B. 6x^2 + 10 C. 6x^2 - 6x - 6 D. -4x^2 - 6x - 6 What is (9x^2-2x+3) - 4x^2 + 7x-5)? A. -5x^2 + 5x - 2 B. 5x^2 - 9x + 8 C. 5x^2 + 5x - 2 D. -5x^2 + 9x - 8

OpenStudy (anonymous):

Its MULTIPLE choice BTW.

OpenStudy (anonymous):

number 1 is C

OpenStudy (anonymous):

Ok!

OpenStudy (zpupster):

remove parentheses and combine like terms x^2 + 5x^2-3x-3x+2 -8

OpenStudy (zpupster):

9x^2- 4x^2 =? -2x+ 7x=? +3 -5=?

OpenStudy (anonymous):

1. 5x^2

OpenStudy (anonymous):

2. 5x 3. -2

OpenStudy (zpupster):

yes

OpenStudy (anonymous):

so it would be C

OpenStudy (zpupster):

good job!!

OpenStudy (anonymous):

Thank you! :)